# Predicting the Reliability of an Image Classifier under Image Distortion

### v1

@@ Abstract
In image classification tasks, deep learning models are vulnerable to image distortions i.e. their accuracy significantly drops if the input images are distorted. An image-classifier is considered “reliable” if its accuracy on distorted images is above a user-specified threshold. For a quality control purpose, it is important to predict if the image-classifier is unreliable/reliable under a distortion level. In other words, we want to predict whether a distortion level makes the image-classifier “non-reliable” or “reliable”. Our solution is to construct a training set consisting of distortion levels along with their “non-reliable” or “reliable” labels, and train a machine learning predictive model (called distortion-classifier) to classify unseen distortion levels. However, learning an effective distortion-classifier is a challenging problem as the training set is highly imbalanced. To address this problem, we propose two Gaussian process based methods to rebalance the training set. We conduct extensive experiments to show that our method significantly outperforms several baselines on six popular image datasets.
@@ 1 Introduction
Although current imbalance handling methods can rebalance the training data, they often suffer from generating false positive samples, leading to a sub-optimal training set for the distortion-classifier $S$. In this paper, we improve the training set ${\cal R}$ of $S$ with two contributions: (1) a Gaussian process (GP) based sampling technique to create a training set ${\cal R}$ with a higher fraction of real positive samples and (2) a GP-based imbalance handling technique to further rebalance ${\cal R}$ by generating more synthetic positive samples.
@@
GP-based sampling: we consider the mapping from a distortion level $c$ to the model’s accuracy on the set of distorted images ${\cal D}^{\prime}_{c}$ as a black-box, expensive function $f:{\cal C}\rightarrow[0,1]$. The function $f$ is black-box as we do not know its expression, and $f$ is expensive as we have to measure the model’s accuracy over all distorted images in ${\cal D}^{\prime}_{c}$. We approximate $f$ using a GP [5] that is a popular method to model black-box, expensive functions. We use $f$’s predictive distribution to design an acquisition function to search for distortion levels that have a high chance to be a positive sample. We update the GP with new samples, and repeat the sampling process until the sampling budget $I$ is depleted. Finally, we obtain a training set =${\cal R}=\{c_{t},\mathbb{I}_{f(c_{t})\geq h}\}_{t=1}^{I}$.
GP-based imbalance handling: we use SMOTE on ${\cal R}$ to generate synthetic positive samples. But, SMOTE often generates many false positive samples [6]. To solve this problem, we assign an uncertainty score for each synthetic positive sample via the variance function of a GP. We filter out synthetic samples whose uncertainty scores are high. This helps to reduce the false positive rate of SMOTE.
@@
    We propose a GP-based imbalance handling method to reduce the false positive rate when generating synthetic positive samples.
@@
The remainder of the paper is organized as follows. In Section 2, related works on image distortion, model reliability, and imbalance classification are reviewed. Our main contributions are presented in Section 3, where we describe two GP-based methods for addressing class imbalance. Experimental results are discussed in Sections 4 while conclusions and future works are represented in Section 5.
@@ 2 Related Work
Imbalance classification. As the problem of classification with imbalanced data has been studied for many years, the imbalance classification has a rich literature. Most existing methods are based on SMOTE (Synthetic Minority Oversampling Technology) [2], where the synthetic minority samples are generated by linearly combining two real minority samples. Several variants have been developed to address SMOTE weaknesses such as outlier and noisy [18, 19, 20, 21, 22]. Other approaches for rebalancing data are under-sampling techniques [3], ensemble methods [23, 24], and generative models [25, 26, 27].
@@ 3.2 Proposed method
We improve the quality of the training set ${\cal R}$ by proposing two new approaches. First, instead of using random sampling method, we propose a GP-based sampling method to sample $c_{i}$ to construct ${\cal R}$. Second, we further rebalance ${\cal R}$ using a novel GP-based imbalance handling technique.
@@ 3.2.1 GP-based sampling
3.2.2 GP-based imbalance handling
We further rebalance the training set ${\cal R}$ by generating synthetic positive samples. Existing over-sampling methods such as SMOTE [2] and its variants [29] suffer from a high false positive rate. To address this problem, we propose a novel method combining SMOTE and GP. We call it SMOTE-GP.
We use SMOTE on ${\cal R}$ to generate synthetic positive samples. Given a real positive sample $c_{i}$, a new synthetic positive sample will be $\hat{c}_{i}=c_{i}+\epsilon\times(c_{j}-c_{i})$, where $c_{j}$ is another real positive sample and $\epsilon\in(0,1)$ is a random number. We indicate the set of synthetic positive samples as $\hat{{\cal R}}^{+}$. However, SMOTE tends to generate false positive samples as the line connecting two real positive samples crosses the negative region, as shown in Figure 3(a). Our goal is to reject such false positive samples.
Figure 3: SMOTE (a) vs. our method SMOTE-GP (b). The black curve indicates the mapping function $f(c)$ (Section 3.1). The red line indicates the threshold $h$ to determine negative samples (blue dots) and positive samples (orange dots). SMOTE generates a synthetic positive sample (green dot) along the line connecting two positive samples, which is false positive (a). SMOTE-GP computes an uncertainty score for the synthetic positive sample via a GP variance function. As the uncertainty score is high, SMOTE-GP rejects this synthetic sample (b).
When SMOTE generates a new synthetic positive sample $\hat{c}_{i}$, it simply assigns $\hat{c}_{i}$ to label $1$. But, SMOTE does not provide any confidence estimation for its assignment. However, if we had such a confidence measure, we could reject the synthetic sample whose confidence is low (i.e. its uncertainty is high).
To measure the uncertainty of a SMOTE assignment, we use the variance function of a GP. First, we retrieve the set of real positive samples in ${\cal R}$, we indicate this set as ${\cal R}^{+}$. Second, we train the GP using real positive samples in ${\cal R}^{+}$ along with their function values. As the GP is trained with only real positive samples, it can approximate the generation process of SMOTE. Then, we compute an uncertainty score $u_{\hat{c}}$ for each synthetic positive sample $\hat{c}\in\hat{{\cal R}}^{+}$ (recall $\hat{{\cal R}}^{+}$ is the set of synthetic positive samples generated by SMOTE):
$u_{\hat{c}}$ $=\sigma^{2}(\hat{c})$
$=k(\hat{c},\hat{c})-\boldsymbol{k}^{\mathsf{T}}K^{-1}\boldsymbol{k},$ (5)
where $\sigma^{2}(\hat{c})$ is the variance function of the GP, $\boldsymbol{k}$ is a vector with its $i$-th element being $k(c_{i},\hat{c})$ and $K$ is a matrix of size $|{\cal R}^{+}|\times|{\cal R}^{+}|$ with its $(i,j)$-th element being $k(c_{i},c_{j})$.
Finally, as $u_{\hat{c}}$ measures the uncertainty of the synthetic positive sample $\hat{c}$, if $u_{\hat{c}}$ is smaller than a threshold $\upsilon$, we keep $\hat{c}$. Otherwise, we discard $\hat{c}$. The procedure of our SMOTE-GP is shown in Algorithm 1.
Input: Imbalanced training set ${\cal R}$ and uncertainty threshold $\upsilon$
1
Output: Rebalanced training set ${\cal R}^{*}$
2
3begin
4
5      ${\cal R}^{*}={\cal R}$
6      generate synthetic positive samples ^$\hat{{\cal R}}^{+}=\text{SMOTE}({\cal R})$
7      train a GP with real positive samples +=${\cal R}^{+}=\{(c_{i},f(c_{i}))\mid c_{i}\in{\cal R}\wedge\text{label}(c_{i})=1\}$
8      for each $\hat{c}_{i}\in\hat{{\cal R}}^{+}$  do
9
10            compute its uncertainty score $u_{\hat{c}_{i}}$ using Equation (5)
11            if $u_{\hat{c}_{i}}\leq$$\upsilon$ then
12
13                  ${\cal R}^{*}={\cal R}^{*}\cup\hat{c}_{i}$
14             end if
15
16       end for
17
18 end
19
Algorithm 1 Our SMOTE-GP algorithm.
@@ 4.1 Experiment settings
Figure 4: Steps involved in the training and test phases of a model reliability prediction task. The module $\text{CT}(\{c_{1},...,c_{I}\}\mid T,{\cal D},h)$ to construct the training set ${\cal R}$ is described in Figure 2. There are two differences between our method and the baseline: (1) the GP-based sampling and (2) the SMOTE-GP.
@@
For image-classifiers $T$, we use five pre-trained models from [33] for image datasets MNIST, Fashion, CIFAR-10, CIFAR-100, and Tiny-ImageNet. They achieved similar accuracy as those reported in [34, 35, 36]. We also use the pre-trained ResNet50 model from the Keras libraryhttps://keras.io/api/applications/resnet/#resnet50-function for ImageNettehttps://www.tensorflow.org/datasets/catalog/imagenette. As pointed out by [12, 1], we expect that these image-classifiers will reduce their performance when evaluated on distorted images.
@@
Imbalance handling method. As the training set ${\cal R}$ is highly imbalanced, we rebalance it before training the distortion-classifier $S$. We use SMOTE-GP and compare it with SOTA imbalance handling methods, including Cost-sensitive learning [37], under-sampling method NearMiss [3], over-sampling methods SMOTE [2] and AdaSyn [38], ensemble methods SPE [23] and MESA [24], and generative models GAN [27], VAE [25], CTGAN, and TVAE [26].
As our method is based on SMOTE, we also compare it with SMOTE variants, including SMOTE-Borderline [20], SMOTE-SVM [21], SMOTE-ENN [19], SMOTE-TOMEK [18], and SMOTE-WB [22]. To be fair, we use the source codes released by the authors or implemented in well-known public libraries. The details are provided in B.
@@ 4.3 Comparison of imbalance handling methods
We compare our imbalance handling method SMOTE-GP with current state-of-the-art imbalance handling methods. Our method has two versions: (1) SMOTE-GP combined with the random sampling and (2) SMOTE-GP combined with our GP-based sampling. We use the uncertainty threshold $\upsilon=0.05$ for CIFAR-10 and $\upsilon=0.005$ for other datasets.
Table 3 shows that our SMOTE-GP combined with our GP-based sampling is the best method and significantly outperforms other methods. Its improvements are around 5% on MNIST, 9% on Fashion, 3% on CIFAR-10, 8% on CIFAR-100, 6% on Tiny-ImageNet, and 14% on ImageNette.
Table 3: F1-scores of our method SMOTE-GP and other imbalance handling methods. Datasets include M: MNIST, F: Fashion, C10: CIFAR-10, C100: CIFAR-100, T-IN: Tiny-ImageNet, and IN: ImageNette.
@@
Ours Random SMOTE-GP 0.6356 0.4611 0.7433 0.6440 0.5856 0.7929
GP-based SMOTE 0.6361 0.5327 0.7562 0.6525 0.5952 0.7988
GP-based SMOTE-GP 0.6635 0.5467 0.7616 0.6780 0.6381 0.8559
When using the random sampling, our SMOTE-GP is still better than other methods by 1-8%. Among imbalance handling baselines, SMOTE often achieves the best results. When SMOTE is combined with our GP-based sampling, its performance is improved significantly. This shows that our GP-based sampling is better than the random sampling.
@@
Comparison with SMOTE variants. We also compare our SMOTE-GP with imbalance handling methods based on SMOTE in Table 4. Our method is the best method while other SMOTE-based methods perform similarly.
Table 4: F1-scores of our method SMOTE-GP and SMOTE variants.
@@ 4.4 Ablation study
Uncertainty threshold $\upsilon$. Our SMOTE-GP uses the uncertainty threshold $\upsilon$ to filter out false positive synthetic samples. We investigate how different values for $\upsilon$ affect our performance.
Figure 7 shows that our SMOTE-GP is always better than SMOTE with a large range of $\upsilon$ values. When $\upsilon$ is too small (i.e. $\upsilon<0.01$), it may drop its F1-score as most of synthetic positive samples are filtered out. When $\upsilon$ is too large (i.e. $\upsilon>0.05$), it may also reduce its F1-score since many false positive samples are introduced.
Figure 7: Our F1-score vs. the uncertainty threshold $\upsilon$.
@@
Figure 8 shows that both methods improve as the number of sampling queries increase as expected. More queries result in more training data and more chance to get positive samples. However, our SMOTE-GP is always better than SMOTE by a large margin.
@@
Visualization. For a quantitative evaluation, we use t-SNE [39] to visualize the synthetic positive samples generated by each method. From Figure 10, SMOTE and its variants generate noisy synthetic samples in two situations. Only our SMOTE-GP avoids these problems.
Figure 10: Visualization of original and synthetic samples on CIFAR-10. Blue and orange dots are real negative and positive samples while green dots are synthetic positive samples generated by each method. Compared to the original data (the top left figure), SMOTE variants suffer from two problems. First, they generate suspicious positive samples in the red circle although there is no original data in this region. Only SMOTE-SVM and our method SMOTE-GP overcome this problem. Second, they generate noisy (most likely wrong) positive samples in the blue square although there is only negative sample in this region. Only SMOTE-WB and our method SMOTE-GP can avoid this issue. In summary, only our SMOTE-GP can overcome two cases where of SMOTE and its variants generate incorrect synthetic samples.
@@ 5 Conclusion
Predicting model reliability is an important task in the quality control process. In this paper, we solve this task in the context of image distortion i.e. we predict if an image-classifier is unreliable/reliable under a distortion level. We form this task as a binary classification process with three main steps: (1) construct a training set, (2) rebalance the training set, and (3) train a distortion-classifier. As the training set is highly imbalanced, we propose two methods to handle the imbalance: (1) a GP-based sampling and (2) SMOTE-GP.
In the GP-based sampling, we approximate the black-box function mapping from a distortion level to the model’s accuracy on distorted images using GP. We then leverage the GP’s mean and variance to form our sampling process.
In the SMOTE-GP method, we compute an uncertainty score for each synthetic positive sample. We then filter out ones whose uncertainty scores are higher than a threshold.
We demonstrate the benefits of our method on six image datasets, where it greatly outperforms other baselines.

### v2

@@ Abstract
In image classification tasks, deep learning models are vulnerable to image distortions i.e. their accuracy significantly drops if the input images are distorted. An image-classifier is considered “reliable” if its accuracy on distorted images is above a user-specified threshold. For a quality control purpose, it is important to predict if the image-classifier is unreliable/reliable under a distortion level. In other words, we want to predict whether a distortion level makes the image-classifier “non-reliable” or “reliable”. Our solution is to construct a training set consisting of distortion levels along with their “non-reliable” or “reliable” labels, and train a machine learning predictive model (called distortion-classifier) to classify unseen distortion levels. However, learning an effective distortion-classifier is a challenging problem as the training set is highly imbalanced. To address this problem, we propose a Gaussian process based method to rebalance the training set. We conduct extensive experiments to show that our method significantly outperforms several baselines on six popular image datasets.
@@ 1 Introduction
Although current imbalance handling methods can rebalance the training data, they often do not generate diverse synthetic positive samples due to a small number of real positive samples in ${\cal R}$. This leads to a sub-optimal training set for the distortion-classifier $S$. In this paper, we improve the training set ${\cal R}$ of $S$ with a Gaussian process (GP) based sampling technique to create a training set ${\cal R}$ with a higher fraction of real positive samples.
@@
GP-based sampling: we consider the mapping from a distortion level $c$ to the model’s accuracy on the set of distorted images ${\cal D}^{\prime}_{c}$ as a black-box, expensive function $f:{\cal C}\rightarrow[0,1]$. The function $f$ is black-box as we do not know its expression, and $f$ is expensive as we have to measure the model’s accuracy over all distorted images in ${\cal D}^{\prime}_{c}$. We approximate $f$ using a GP [33, 27] that is a popular method to model black-box, expensive functions. We use $f$’s predictive distribution to design an acquisition function to search for distortion levels that have a high chance to be a positive sample. We update the GP with new samples, and repeat the sampling process until the sampling budget $I$ is depleted. Finally, we obtain a training set =${\cal R}=\{c_{t},\mathbb{I}_{f(c_{t})\geq h}\}_{t=1}^{I}$.
@@
The remainder of the paper is organized as follows. In Section 2, related works on image distortion, model reliability, and imbalance classification are reviewed. Our main contributions are presented in Section 3, where we describe a GP-based sampling method. Experimental results are discussed in Sections 4 while conclusions and future works are represented in Section 5.
@@ 2 Related Work
Imbalance classification. As the problem of classification with imbalanced data has been studied for many years, the imbalance classification has a rich literature. Most existing methods are based on SMOTE (Synthetic Minority Oversampling Technology) [7], where the synthetic minority samples are generated by linearly combining two real minority samples. Several SMOTE variants have been developed to address SMOTE weaknesses such as outlier and noisy [2, 3, 13, 28, 30]. Other approaches for rebalancing data are under-sampling techniques [22], ensemble methods [20, 21], and generative models [18, 40, 10].
@@ 3.2 Proposed method
We improve the quality of the training set ${\cal R}$ by proposing a new sampling approach. Instead of using random sampling method, we propose a GP-based sampling method to sample $c_{i}$ to construct ${\cal R}$.
@@ 4.1 Experiment settings
Figure 3: Steps involved in the training and test phases of a model reliability prediction task. The module $\text{CT}(\{c_{1},...,c_{I}\}\mid T,{\cal D},h)$ to construct the training set ${\cal R}$ is described in Figure 2. There is one difference between our method and the baseline: the GP-based sampling.
@@
For image-classifiers $T$, we use five pre-trained models from [25] for image datasets MNIST, Fashion, CIFAR-10, CIFAR-100, and Tiny-ImageNet. They achieved similar accuracy as those reported in [37, 38, 26, 4]. We also use the pre-trained ResNet50 model from the Keras library***https://keras.io/api/applications/resnet/#resnet50-function for ImageNettehttps://www.tensorflow.org/datasets/catalog/imagenette. As pointed out by [15, 19], we expect that these image-classifiers will reduce their performance when evaluated on distorted images.
@@
Imbalance handling method. As the training set ${\cal R}$ is highly imbalanced, we rebalance it before training the distortion-classifier $S$. We use SMOTE [7] and compare it with SOTA imbalance handling methods, including Cost-sensitive learning [35], under-sampling method NearMiss [22], over-sampling methods AdaSyn [14], ensemble methods SPE [20] and MESA [21], and generative models GAN [10], VAE [18], CTGAN, and TVAE [40].
As our method is SMOTE, we also compare it with SMOTE variants, including SMOTE-Borderline [13], SMOTE-SVM [28], SMOTE-ENN [3], SMOTE-TOMEK [2], and SMOTE-WB [30]. To be fair, we use the source codes released by the authors or implemented in well-known public libraries. The details are provided in B.
@@ 4.3 Comparison of imbalance handling methods
We compare our imbalance handling method GP-based sampling + SMOTE (we call it GPS-SMOTE) with current state-of-the-art imbalance handling methods.
Table 3 shows that our GPS-SMOTE is the best method and significantly outperforms other methods. Its improvements are around 2% on MNIST, 8% on Fashion, 2% on CIFAR-10, 6% on CIFAR-100, 2% on Tiny-ImageNet, and 8% on ImageNette.
Table 3: F1-scores of our method GPS-SMOTE and other imbalance handling methods. Datasets include M: MNIST, F: Fashion, C10: CIFAR-10, C100: CIFAR-100, T-IN: Tiny-ImageNet, and IN: ImageNette.
@@
GPS-SMOTE (Ours) GP-based SMOTE 0.6361 0.5327 0.7562 0.6525 0.5952 0.7988
Among imbalance handling baselines, SMOTE often achieves the best results. When SMOTE is combined with our GP-based sampling, its performance is improved significantly. This shows that our GP-based sampling is better than the random sampling.
@@
Comparison with SMOTE variants. We also compare our GPS-SMOTE with imbalance handling methods based on SMOTE in Table 4. Our method is the best method while other SMOTE-based methods perform similarly.
Table 4: F1-scores of our method GPS-SMOTE and SMOTE variants.
@@ 4.4 Ablation study
Figure 6 shows that both methods improve as the number of sampling queries increase as expected. More queries result in more training data and more chance to get positive samples. However, our GPS-SMOTE is always better than SMOTE by a large margin.
@@ 5 Conclusion
Predicting model reliability is an important task in the quality control process. In this paper, we solve this task in the context of image distortion i.e. we predict if an image-classifier is unreliable/reliable under a distortion level. We form this task as a binary classification process with three main steps: (1) construct a training set, (2) rebalance the training set, and (3) train a distortion-classifier. As the training set is highly imbalanced, we propose a method to handle the class-imbalance: a GP-based sampling. Namely, we approximate the black-box function mapping from a distortion level to the model’s accuracy on distorted images using GP. We then leverage the GP’s mean and variance to form our sampling process. We demonstrate the benefits of our method on six image datasets, where it greatly outperforms other baselines.
Future works. We will investigate and verify our method for other ML and deep learning models including large language models [6], generative models [10], and tabular models [12, 24].